# Empirical Analysis for Unsupervised Universal Dependency Parse Tree Aggregation

### v1

@@ 5.3 Baseline Methods
The three aggregation frameworks are compared with top two ensemble methods, including HIT-SCIR Che et al. (2018) and LATTICE Lim et al. (2018), top two non-ensemble methods, including TurkuNLP Kanerva et al. (2018) and UDPipe Future Straka (2018) from CoNLL 2018 shared task, and two LLM-based methods, UAdapter Üstün et al. (2020) and MLPSBM Xu et al. (2022). Following their respective papers’ settings, we use the BERT-multilingual-cased encoder for both UAdapter and MLPSBM. MLPSBM is tested only for high-resource languages such as Bulgarian, Catalan, Czech, German, English, Spanish, Italian, Dutch, Norwegian, Romanian, and Russian. For reference, we also present the performance of the best parser among these top 9 chosen parsers (BEST) and the average performance of the top 9 chosen parsers (Average) for each test treebank evaluated using ground truth annotations.
@@ A.4 Full Results
Full results [^6] We will publish the code upon acceptance for $19$ high-resource language treebanks are shown in Table 3 and the full results for $52$ low-resource language treebanks are shown in Tables 4 and 5. The table is ordered alphabetically concerning the name of the treebank. From the results in Table 3, taking treebank $bg\_btb$ as an example, we can observe that MLPSBM, the LLM-based parser, achieved the best results among the baselines and CIM outperformed all the baselines for this treebank. Similar observation can be made for results in Table 4 for $af\_afribooms$ treebank where UAdapter, another LLM-based parser, achieved the best results among the baselines and CIM outperformed all the baselines for this treebank too. The examples confirm that CIM is the most suitable DTS aggregation framework that can properly estimate parser quality and outperform state-of-the-art baselines.
@@ A.5 Comparison Study
Figures 1, 2, and 3 depict the histograms of comparisons between CIM with ensemble, non-ensemble, and LLM-based baselines for high-resource language treebanks. A positive difference means that CIM outperforms. From figure 1, we can observe that CIM outperforms ensemble methods, including HIT-SCIR and LATTICE. From figure 2, we can observe that CIM outperforms non-ensemble methods, including TurkuNLP and UDPipe Future. From figure 3, we can observe that CIM outperforms LLM-based baselines UAdapter and MLPSBM. These results support our claim that CIM can correct the mistakes of individual parsers — even the best ones — and outperform them. These results support our claim that CIM is the most suitable DTS aggregation framework among the compared aggregation frameworks in this study.
Table 3: Full results of MST, CRH, CIM compared with the baselines for $19$ high-resource language treebanks. The table is ordered in alphabetical order with respect to the name of the treebank. The columns E1 and E2 refer to ensemble baselines HIT-SCIR Che et al. (2018), and LATTICE Lim et al. (2018), respectively. The columns NE1, NE2 refer to non-ensemble baselines TurkuNLP Kanerva et al. (2018) and UDPipe Future Straka (2018), respectively. The columns L1 and L2 refer to LLM-based parsers UAdapter Üstün et al. (2020) and MLPSBM Xu et al. (2022), respectively. The column Avg. refers to the baseline Average.
@@
Table 4: Full results of MST, CRH, CIM compared with the baselines for $52$ low-resource language treebanks. The table is ordered in alphabetical order with respect to the name of the treebank. The columns E1 and E2 refer to ensemble baselines HIT-SCIR Che et al. (2018), and LATTICE Lim et al. (2018), respectively. The columns NE1, NE2 refer to non-ensemble baselines TurkuNLP Kanerva et al. (2018) and UDPipe Future Straka (2018), respectively. The column L1 refers to LLM-based parser UAdapter Üstün et al. (2020). The column Avg. refers to the baseline Average.
@@
Table 5: Full results of MST, CRH, CIM compared with the baselines for $52$ low-resource language treebanks (Contd.). The table is ordered in alphabetical order with respect to the name of the treebank. The columns E1 and E2 refer to ensemble baselines HIT-SCIR Che et al. (2018), and LATTICE Lim et al. (2018), respectively. The columns NE1, NE2 refer to non-ensemble baselines TurkuNLP Kanerva et al. (2018) and UDPipe Future Straka (2018), respectively. The column L1 refers to LLM-based parser UAdapter Üstün et al. (2020). The column Avg. refers to the baseline Average.

### v2

@@ 5.3 Baseline Methods
The three aggregation frameworks are compared with top two ensemble methods, including HIT-SCIR Che et al. (2018) and LATTICE Lim et al. (2018), top two non-ensemble methods, including TurkuNLP Kanerva et al. (2018) and UDPipe Future Straka (2018) from CoNLL 2018 shared task, and two LLM-based methods, UDapter Üstün et al. (2020) and MLPSBM Xu et al. (2022). Following their respective papers’ settings, we use the BERT-multilingual-cased encoder for both UDapter and MLPSBM. MLPSBM is tested only for high-resource languages such as Bulgarian, Catalan, Czech, German, English, Spanish, Italian, Dutch, Norwegian, Romanian, and Russian. For reference, we also present the performance of the best parser among these top 9 chosen parsers (BEST) and the average performance of the top 9 chosen parsers (Average) for each test treebank evaluated using ground truth annotations.
@@ A.4 Full Results
Full results [^6] We will publish the code upon acceptance for $19$ high-resource language treebanks are shown in Table 3 and the full results for $52$ low-resource language treebanks are shown in Tables 4 and 5. The table is ordered alphabetically concerning the name of the treebank. From the results in Table 3, taking treebank $bg\_btb$ as an example, we can observe that MLPSBM, the LLM-based parser, achieved the best results among the baselines and CIM outperformed all the baselines for this treebank. Similar observation can be made for results in Table 4 for $af\_afribooms$ treebank where UDapter, another LLM-based parser, achieved the best results among the baselines and CIM outperformed all the baselines for this treebank too. The examples confirm that CIM is the most suitable DTS aggregation framework that can properly estimate parser quality and outperform state-of-the-art baselines.
@@ A.5 Comparison Study
Figures 1, 2, and 3 depict the histograms of comparisons between CIM with ensemble, non-ensemble, and LLM-based baselines for high-resource language treebanks. A positive difference means that CIM outperforms. From figure 1, we can observe that CIM outperforms ensemble methods, including HIT-SCIR and LATTICE. From figure 2, we can observe that CIM outperforms non-ensemble methods, including TurkuNLP and UDPipe Future. From figure 3, we can observe that CIM outperforms LLM-based baselines UDapter and MLPSBM. These results support our claim that CIM can correct the mistakes of individual parsers — even the best ones — and outperform them. These results support our claim that CIM is the most suitable DTS aggregation framework among the compared aggregation frameworks in this study.
Table 3: Full results of MST, CRH, CIM compared with the baselines for $19$ high-resource language treebanks. The table is ordered in alphabetical order with respect to the name of the treebank. The columns E1 and E2 refer to ensemble baselines HIT-SCIR Che et al. (2018), and LATTICE Lim et al. (2018), respectively. The columns NE1, NE2 refer to non-ensemble baselines TurkuNLP Kanerva et al. (2018) and UDPipe Future Straka (2018), respectively. The columns L1 and L2 refer to LLM-based parsers UDapter Üstün et al. (2020) and MLPSBM Xu et al. (2022), respectively. The column Avg. refers to the baseline Average.
@@
Table 4: Full results of MST, CRH, CIM compared with the baselines for $52$ low-resource language treebanks. The table is ordered in alphabetical order with respect to the name of the treebank. The columns E1 and E2 refer to ensemble baselines HIT-SCIR Che et al. (2018), and LATTICE Lim et al. (2018), respectively. The columns NE1, NE2 refer to non-ensemble baselines TurkuNLP Kanerva et al. (2018) and UDPipe Future Straka (2018), respectively. The column L1 refers to LLM-based parser UDapter Üstün et al. (2020). The column Avg. refers to the baseline Average.
@@
Table 5: Full results of MST, CRH, CIM compared with the baselines for $52$ low-resource language treebanks (Contd.). The table is ordered in alphabetical order with respect to the name of the treebank. The columns E1 and E2 refer to ensemble baselines HIT-SCIR Che et al. (2018), and LATTICE Lim et al. (2018), respectively. The columns NE1, NE2 refer to non-ensemble baselines TurkuNLP Kanerva et al. (2018) and UDPipe Future Straka (2018), respectively. The column L1 refers to LLM-based parser UDapter Üstün et al. (2020). The column Avg. refers to the baseline Average.